The target diameter of bolts from a production line is 10mm. Historically the bolt diameters are normally distributed with a standard deviation of 0.15mm. To monitor this process periodically an engineer takes a random sample of 6 measurements. Let µ be the true average bolt diameter.

Ho:u=10 versus Ha:+10


The rejection region is:   >10+c or ī<10-C
Find the minimum value of c that yields a test with significance 0.01?

zα 1.282 1.645 1.960 2.326 2.576 3.090
α(tail area) 0.1 0.05 0.025 0.01 0.005 0.001
%ile 90 95 97.5 99 99.5 99.9


Your answer can be rounded to four decimal digit accuracy when entered.
Answer #1

At alpha = 0.01, the critical values are +/- z0.005 = +/- 2.576

z > 2.576

or, (エー)/(ulo) > 2.576

or, (\bar x - 10)/(0.15/\sqrt 6) > 2.576

or, \bar x> 2.576 * 0.15/\sqrt 6 + 10

or, \bar x> 10.1577

\bar x> 10 + c

or, 10.1577 = 10 + c

or, 0.1577 = c

z < -2.576

or, (エー)/(ulo) < -2.576

or, (\bar x - 10)/(0.15/\sqrt 6) < -2.576

or, \bar x< -2.576 * 0.15/\sqrt 6 + 10

or, \bar x< 9.8423

\bar x< 10 - c

or, 9.8423 = 10 - c

or, c = 0.1577

So the minimum value of c is 0.1577.